# arrows,decorations.markings,shapes,matrix,positioning, calc, fit,decorations.pathreplacing,angles,quotes,backgrounds

### v1

@@ 5.1. Metrics
We evaluate study outcomes based on participants’ perceptions and decisions. We consider two fairness perception metrics based on questions in the post-phase surveys, and we consider the decisions made by the human-AI teams based on one fairness measure—demographic parity—and three decision quality metrics: accuracy, false negative rate, and false positive rate. We measure all metrics in the two task phases across conditions. We count both “Likely Complete” and “Definitely Complete” as “Complete” and similarly for “Late”. We count “Neutral” as agreement with the system prediction.
@@ 5.1.1. Decision-Making Fairness Measure
We employ demographic parity (Feldman et al., 2015, i.a.) as a measure of fairness in decision-making, which captures the independence between protected characteristics and prediction. There are other measures of fairness (Narayanan, 2018), however, not all definitions can be simultaneously satisfied (Chouldechova, 2017; Kleinberg et al., 2017). Demographic parity has been found to be more understandable to laypeople and better capture their perception of fairness than competing metrics (Saha et al., 2020; Srivastava et al., 2019). We can calculate the demographic parity for human-AI teams in task phases 1 and 2 across conditions as follows. Parity =
E[^Yi,j=1 ∣Gender=female]E[^Yi,j=1 ∣Gender=male], where $\hat{Y}_{i,j}$ is the predicted decision for the applicant by the participant $j$. We obtain one demographic parity score in this way for each participant’s decisions in each phase. A parity close to $1$ means an equal acceptance rate. As the acceptance rate for the advantaged group increases over the disadvantaged group, parity becomes closer to 00. If the acceptance rate of the disadvantaged group increases above the advantaged group, then the parity can increase above $1$. A parity of less than $\frac{4}{5}$ is considered “evidence of adverse impact” under US Anti-Discrimination law (Equal Employment Opportunity Commission, 1978). In our model bias disclosure, we tell the participants about this $80\%$ rule and that the model failed this test in phase 1, that is, the demographic parity of the model is below $80\%$ (2(a)).
@@ 5.1.3. Decision Quality Measures
[…]
@@ 5.2. Statistical Analyses
To answer our research questions (§3), we perform separate multi-way ANOVA tests for different treatments (explanations, disclosures without explanations, and disclosures with explanations) for both protected and proxy conditions. For each statistical test, we construct a linear model with a fixed effect term for each independent treatment variable and one fixed effect term representing the participant’s dispositional trust, which is calculated by averaging the scores from the pre-study trust survey. Additionally, in the within study comparisons (that is, moving from phase 1 to phase 2), we include the participant ID as a random effect. The independent treatment variables are determined by the factors that vary between the effect of interest. For instance, to estimate the effect of explanation (that is, the vertical arrow in §5.1.3), the treatment variable is the presence of explanations. For the effect of disclosure (that is, the horizontal arrows in §5.1.3), the treatment variables are: (1) whether only bias disclosure has been shown (i.e., is this a phase 2 measurement with no correlation disclosure), and (2) whether full bias and correlation disclosure has been shown. For the effect of adding both explanations and disclosures (that is, the diagonal in §5.1.3 going from without explanation and disclosures in phase 1 to with explanation and disclosures in phase 2), we include all three treatment variables. In each ANOVA test, we consider the data from the relevant sections. For instance, to estimate the effect of explanations alone in the case of direct bias through a protected feature, we only consider the data in phase 1 of the “protected” conditions (left vertical section in §5.1.3), and similarly for “proxy” conditions. We fit a separate model for each fairness perception and decision-making metric as the dependent variable for each of the above effects. Although the ratio of “Complete” and “Late” model decisions shown to each participant is kept the same in each phase across conditions (leading to a constant AI-only parity of $\sim$0.670.670.670.67), the model accuracy, FPR, and FNR varies across conditions. To account for this variation, we subtract the model score from the score of the human-AI team. Similarly, when considering learned trust measures, we also adjust for dispositional trust in AI by subtracting the participant’s answer to the pre-study survey for the corresponding question. In addition to our key metrics detailed in §5.1, we also study the effect of the treatments on learned trust. We follow the same procedure as before, but with the learned trust measures as the dependent variable in the ANOVA test. In this case, however, we adjust participants’ post-phase 1 or post-phase 2 survey responses based on their baseline responses, and we do not include the overall dispositional trust term. Lastly, we perform additional ANOVA tests considering the difference between dispositional and learned trust. For this, we consider the participants’ trust ratings in the pre-study survey and surveys after phase 1 or phase 2 as the dependent variable. We fit two linear models, one for phase 1 and 2 each, testing whether the phase has a fixed effect on the participants’ trust ratings in different conditions (with the participant added as a random effect). We perform Benjamini-Hochberg correction to avoid multiple testing effect with a false discovery threshold of $0.05$ (Benjamini and Hochberg, 1995). This leads to a significance threshold of $0.0175$ for the reported results.
@@ 6. Quantitative Results
In this section, we report our findings on the effects of different interventions on the decision-making and fairness perception metrics. We first discuss the primary effects detailed in §5.1.3: the effect of explanations (§6.1), the effect of disclosures without explanations (§6.2) and the effect of disclosures with explanations (§6.3). Next, we consider the effect of the joint intervention of adding both explanations and disclosures to assess their aggregate benefits, if any (§6.4). Lastly, we discuss the effect of dispositional trust on the decision-making and fairness perception, the effect of different interventions on participants’ trust, and the differences in participants’ dispositional trust vs trust in our system in §6.5 (full results in Appendix B).
[…]
@@ 6.1. Effect of Explanations Alone
First, we consider the effects of explanations alone by comparing between the first phase of with and without explanations conditions with either type of bias. In the case of direct bias through a protected feature, we find that explanations alone have a significant effect on all metrics; however, the direction of the effect is not consistent. Explanations alone significantly improve participants’ ability to recognize unfairness (§6). Surprisingly, despite participants being more able to recognize that the model is unfair when shown explanations, when considering decisions instead of perceptions, we see that explanations significantly decrease gender parity (§6). Looking closer, we find that explanations lead to a significantly lower acceptance rate for female applicants, whereas the acceptance rate for male applicants does not change significantly (Table 5 in the Appendix). This decrease in acceptance rates also leads to a significant increase in the FNR and a significant decrease in FPR, with an overall higher accuracy (§6). In the case of indirect bias through a proxy feature, we find that explanations alone significantly reduce gender parity, similar to the case of direct bias. (§6). Analogous to the direct bias, this occurs due to a significant decrease in acceptance of female applicants (Table 5 in the Appendix). Further, explanations also lead to a significant increase in accuracy in the case of indirect bias as well. However, unlike in the case of direct bias, explanations have no significant effect on fairness perceptions in the case of indirect bias (§6). Overall, we find that explanations can help people recognize unfairness in the case of direct bias but not indirect. However, regardless of fairness perceptions, in line with Wang et al. (2023), we find that explanations lead people to accept model biases leading to less fair decisions.
[…]
@@ 6.2. Effect of Disclosures without Explanations
We consider the effects of disclosures without explanations by comparing between phase 1 and phase 2 in without explanations conditions with either type of bias. In the case of both direct and indirect bias, we find that disclosing model bias alone does not have a significant effect on any of the outcome metrics (gender parity, fairness perception, accuracy, FPR, and FNR). However, in the case of indirect bias, when we disclose both the model bias and the relationship between the protected and proxy feature (i.e., that some universities in the study are women’s colleges), participants were significantly more likely to report that the model is unfair or that this unfairness caused them to disagree with the model’s decisions (§6.1). Interestingly, this still does not translate to fairer decisions—as seen in §6.1, the gender parity does not change significantly on disclosing both model bias and correlations in the case of indirect bias. In sum, we find that, interestingly, being explicitly told that the model is biased does not affect participants’ fairness perception of the model decisions (in both direct and indirect bias conditions). But, disclosing both the model bias and the correlation between protected and proxy feature does lead to participants perceiving the model as less fair in the case of indirect bias. However, this is not sufficient to improve the decision-making fairness.
[…]
@@ 6.3. Effect of Disclosures with Explanations
We consider the effects of disclosures with explanations by comparing between phase 1 and phase 2 in explanations conditions with either type of bias. In the case of direct bias through a protected feature, we find that bias disclosure with explanations has no significant effect on fairness perceptions (§6.2). Different from bias disclosure without explanations (§6.2), we find that bias disclosure with explanations significantly increases the acceptance rate for female applicants (participants flip models’ “Late” prediction for female applicants at a much higher rate), with the acceptance rate for the male applicants unchanged (Table 5 in the Appendix). Bias disclosure with explanations also results in a significant increase in FPR and a significant decrease in FNR, leading to an overall insignificant change in accuracy (§6.2). Even despite a higher rate of acceptance for female applicants, the increase in gender parity is not significant (§6.2). This is likely due to the normalizing effect of the acceptance rate of male applicants, which also increases, but not significantly. In the case of indirect bias through a proxy feature, we find that disclosures with explanations have a positive impact on fairness both with respect to perceptions and decision-making. Similarly to what we saw in §6.2, disclosing both model bias and the association between gender and university while including explanations significantly decreases perceived model fairness (§6.2). In the case of fairness rating, this effect is significant even without the correlation portion of the disclosure. Further, as seen in §6.2, disclosing model bias alone, as well as disclosing model bias along with the correlation between protected and proxy feature with explanations leads to a significant increase in gender parity. This stems from a significantly higher acceptance rate for female applicants (Table 5 in the Appendix). The acceptance rate for male applicants remains unchanged. This also results in a higher FPR (significant) and a lower FNR (not significant), with an overall drop in accuracy (significant). In sum, we find that in the case of direct bias, even though bias disclosure with explanations does not improve recognition of model unfairness or improve the decision-making quality or fairness significantly, it does reduce agreement with model’s biased decisions, leading to a significantly higher acceptance rate for female applicants. This is in stark contrast with the effect of explanations alone (§6.1), which improved recognition of fairness, but led to more biased decisions overall. Further, in the case of indirect bias, disclosing the model bias and correlations between protected and proxy feature with explanations significantly increases both recognition of fairness and gender parity in decision-making. We also observe an approximately $1\%$ drop in accuracy in the case of indirect bias after disclosing model bias and correlations with explanations (which might be acceptable in certain cases). Overall, we conclude that neither explanations nor bias or correlation disclosures alone is sufficient. We observe better decision-making fairness outcomes when participants are not only shown the model explanations, but also made aware of the biases underlying them.
@@ 6.4. Effect of Joint Intervention
We have seen that explanations alone decrease decision-making fairness, while disclosures with explanations can, in the case of indirect bias, have the opposite effect. Here, we consider the effect of adding explanations and giving disclosures over including neither (i.e., comparing phase 1 without explanations to phase 2 with explanations). We find that, for decision-making metrics, the effect of joint intervention is never significant (Table 6 in the Appendix). For fairness perception metrics, since the effect of explanations alone and disclosures with explanations pointed in the same direction, as expected, adding both explanations and disclosures also significantly improves recognition of unfairness. In sum, the joint intervention of including both explanations and disclosures (over including neither) helps participants in recognizing model biases but not in correcting them.
@@ 7. Qualitative Results
As described in §4.2, for a selected set of applicants in each phase, we also ask participants to write a free text response after they have made their prediction, with a justification for why they agreed or disagreed with AI (or marked it as “neutral”). In addition to encouraging careful thinking, this also helps us gauge the kinds of reasoning participants employ in their decision-making. As the main goal of our study is to understand how humans interact with AI decisions when the AI is biased, we primarily focus our qualitative analysis on rationales concerning biases. To analyze how participants perceive and use (or discard) the biased feature (gender or university), we consider justifications that directly reference the protected (gender) or proxy feature (university) by using a set of keywords for both. For keywords, we started with an initial set (e.g., “gender”, “female”, “university”) and, based on reading a subset of the justifications, expanded to include spelling variations (e.g., “skool” and “collage”) and other topically relevant words (e.g., abbreviated names of schools). We discuss our qualitative findings on justifications involving the protected feature in §7.1 and justifications involving the proxy feature in §7.2. Lastly, we discuss additional observations indicating over-reliance based on a random sample of justifications in §7.3.

### v2

@@ 5.1. Metrics
We evaluate study outcomes based on participants’ perceptions and decisions. We consider two fairness perception metrics based on questions in the post-phase surveys, and we consider the decisions made by the human-AI teams based on one fairness measure—demographic parity—and three decision quality metrics: accuracy, false negative rate, and false positive rate.
We measure all metrics in the two task phases across conditions. We count both “Likely Complete” and “Definitely Complete” as “Complete” and similarly for “Late”. We count “Neutral” as agreement with the system prediction.
@@ 5.1.1. Decision-Making Fairness Measure
We employ demographic parity (Feldman et al., 2015, i.a.) as a measure of fairness in decision-making, which captures the independence between protected characteristics and prediction. There are other measures of fairness (Narayanan, 2018), however, not all definitions can be simultaneously satisfied (Chouldechova, 2017; Kleinberg et al., 2017). Demographic parity has been found to be more understandable to laypeople and better capture their perception of fairness than competing metrics (Saha et al., 2020; Srivastava et al., 2019). We can calculate the demographic parity for human-AI teams in task phases 1 and 2 across conditions as follows.
$\text{Parity}=\frac{\mathbb{E}[\hat{Y}_{i,j}=1\mid\text{Gender}=\texttt{female}]}{\mathbb{E}[\hat{Y}_{i,j}=1\mid\text{Gender}=\texttt{male}]},$
where $\hat{Y}_{i,j}$ is the predicted decision for the applicant by the participant $j$. We obtain one demographic parity score in this way for each participant’s decisions in each phase.
A parity close to $1$ means an equal acceptance rate. As the acceptance rate for the advantaged group increases over the disadvantaged group, parity becomes closer to 00. If the acceptance rate of the disadvantaged group increases above the advantaged group, then the parity can increase above $1$. A parity of less than $\frac{4}{5}$ is considered “evidence of adverse impact” under US Anti-Discrimination law (Equal Employment Opportunity Commission, 1978). In our model bias disclosure, we tell the participants about this $80\%$ rule and that the model failed this test in phase 1, that is, the demographic parity of the model is below $80\%$ (3(a)).
@@ 5.1.3. Decision Quality Measures
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Fairness Perceptions (§5.1.2)
(b) Decision-Making Fairness (§5.1.1)
(c) Decision-Making Quality (§5.1.3)
Figure 5. Effect of explanations alone on various metrics when bias stems from usage of a protected vs proxy feature. The marks show the average and standard error of the given metric across participants in the given condition.
@@ 5.2. Statistical Analyses
To answer our research questions (§3), we perform separate multi-way ANOVA tests for different treatments (explanations, disclosures without explanations, and disclosures with explanations) for both protected and proxy conditions. For each statistical test, we construct a linear model with a fixed effect term for each independent treatment variable and one fixed effect term representing the participant’s dispositional trust, which is calculated by averaging the scores from the pre-study trust survey. Additionally, in the within study comparisons (that is, moving from phase 1 to phase 2), we include the participant ID as a random effect.
The independent treatment variables are determined by the factors that vary between the effect of interest. For instance, to estimate the effect of explanation (that is, the vertical arrow in Figure 1), the treatment variable is the presence of explanations. For the effect of disclosure (that is, the horizontal arrows in Figure 1), the treatment variables are: (1) whether only bias disclosure has been shown (i.e., is this a phase 2 measurement with no correlation disclosure), and (2) whether full bias and correlation disclosure has been shown. For the effect of adding both explanations and disclosures (that is, the diagonal in Figure 1 going from without explanation and disclosures in phase 1 to with explanation and disclosures in phase 2), we include all three treatment variables.
In each ANOVA test, we consider the data from the relevant sections. For instance, to estimate the effect of explanations alone in the case of direct bias through a protected feature, we only consider the data in phase 1 of the “protected” conditions (left vertical section in Figure 1), and similarly for “proxy” conditions.
We fit a separate model for each fairness perception and decision-making metric as the dependent variable for each of the above effects. Although the ratio of “Complete” and “Late” model decisions shown to each participant is kept the same in each phase across conditions (leading to a constant AI-only parity of $\sim$0.670.670.670.67), the model accuracy, FPR, and FNR vary across conditions. To account for this variation, we subtract the model score from the score of the human-AI team. Similarly, in assessing learned trust measures, we adjust for dispositional trust in AI by subtracting the participant’s response to the corresponding question in the pre-study survey.
In addition to our key metrics detailed in §5.1, we also study the effect of the treatments on learned trust. We follow the same procedure as before but with the learned trust measures as the dependent variable in the ANOVA test. In this case, however, we adjust participants’ post-phase 1 or post-phase 2 survey responses based on their baseline responses, and we do not include the overall dispositional trust term. Lastly, we perform additional ANOVA tests to analyze the difference between dispositional and learned trust. For this, we use the participants’ trust ratings in the pre-study survey and surveys after phase 1 or phase 2 as the dependent variable. We fit two linear models, one for each phase, testing whether the phase has a fixed effect on the participants’ trust ratings in different conditions (with the participant added as a random effect).
We perform Benjamini-Hochberg correction to avoid multiple testing effect with a false discovery threshold of $0.05$ (Benjamini and Hochberg, 1995). This leads to a significance threshold of $0.0175$ for the reported results.
@@ 6. Quantitative Results
In this section, we report our findings on the effects of different interventions on the decision-making and fairness perception metrics. We first discuss the primary effects detailed in Figure 1: the effect of explanations (§6.1), the effect of disclosures without explanations (§6.2) and the effect of disclosures with explanations (§6.3). Next, we examine the effect of the joint intervention of adding both explanations and disclosures (§6.4). Lastly, we discuss the effect of dispositional trust on the decision-making and fairness perception, the effect of different interventions on participants’ trust, and the differences in participants’ dispositional trust vs learned trust in §6.5 (full results in Appendix B).
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Fairness Perceptions (§5.1.2)
(b) Decision-Making Fairness (§5.1.1)
(c) Decision-Making Quality (§5.1.3)
Figure 6. Effect of disclosures without explanations on various metrics when bias stems from usage of a protected vs proxy feature. The marks show the average and standard error of the given metric across participants in the given condition.
@@ 6.1. Effect of Explanations Alone
First, we consider the effects of explanations alone by comparing the first phase of with and without explanations conditions with either type of bias.
In the case of direct bias through a protected feature, we find that explanations alone have a significant effect on all metrics; however, the direction of the effect is not consistent. Explanations alone significantly improve participants’ ability to recognize unfairness (Figure 5). Surprisingly, despite participants being more able to recognize that the model is unfair when shown explanations, when considering decisions instead of perceptions, we see that explanations significantly decrease gender parity (Figure 5). Looking closer, we find that explanations lead to a significantly lower acceptance rate for female applicants, whereas the acceptance rate for male applicants does not change significantly (Table 5 in the Appendix). This decrease in acceptance rates also leads to a significant increase in the FNR and a significant decrease in FPR, with an overall higher accuracy (Figure 5).
In the case of indirect bias through a proxy feature, we find that explanations alone significantly reduce gender parity, similar to the case of direct bias (Figure 5). Analogous to direct bias, this occurs due to a significant decrease in acceptance of female applicants (Table 5 in the Appendix). Further, explanations also lead to a significant increase in accuracy in the case of indirect bias. However, unlike with direct bias, explanations have no significant effect on fairness perceptions in the case of indirect bias (Figure 5).
Overall, we find that explanations can help people recognize unfairness in the case of direct bias but not indirect. This is in line with our intuition that indirect biases are harder for participants to notice. However, regardless of fairness perceptions, in line Wang et al. (2023), we find that explanations lead people to accept model biases leading to less fair decisions. This could be attributed to the presence of explanations assisting humans in rationalizing AI’s unfair predictions rather than challenging them.
@@ 6.2. Effect of Disclosures without Explanations
We consider the effects of disclosures without explanations by comparing between phase 1 and phase 2 in without explanations conditions with either type of bias.
For both direct and indirect bias, we find that disclosing model bias alone does not have a significant effect on any of the outcome metrics (gender parity, fairness perception, accuracy, FPR, and FNR).
However, in the case of indirect bias, when we disclose both the model bias and the relationship between the protected and proxy feature (i.e., that some universities in the study are women’s colleges), participants were significantly more likely to report that the model is unfair or that this unfairness caused them to disagree with the model’s decisions (Figure 6). Interestingly, this still does not translate to fairer decisions—as seen in Figure 6, the gender parity does not change significantly on disclosing both model bias and correlations in the case of indirect bias.
In sum, we find that, interestingly, being explicitly told that the model is biased does not affect participants’ fairness perception of the model decisions (in both direct and indirect bias conditions). In the direct bias condition, this could be because the model is perceived as unfair even pre-disclosures. In the indirect bias condition, this might be because disclosures alone, without explanations, are insufficient for participants to fully acknowledge the bias in the model’s predictions. But, disclosing both the model bias and the correlation between protected and proxy features does lead to participants perceiving the model as less fair in the case of indirect bias. However, this is not sufficient to improve decision-making fairness. This may be because, although disclosures assist participants in recognizing the unfairness of model predictions, they still lack sufficient information to overturn individual predictions without additional guidance on how the model utilizes the correlation between protected and proxy features.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Fairness Perceptions (§5.1.2)
(b) Decision-Making Fairness (§5.1.1)
(c) Decision-Making Quality (§5.1.3)
Figure 7. Effect of disclosures with explanations on various metrics when bias stems from usage of a protected vs proxy feature. The marks show the average and standard error of the given metric across participants in the given condition.
@@ 6.3. Effect of Disclosures with Explanations
We consider the effects of disclosures with explanations by comparing between phase 1 and phase 2 in explanations conditions with either type of bias.
In the case of direct bias through a protected feature, we find that bias disclosure with explanations has no significant effect on fairness perceptions (Figure 7). Unlike with bias disclosure without explanations (§6.2), we find that bias disclosure with explanations significantly increases the acceptance rate for female applicants (participants flip models’ “Late” predictions for female applicants at a much higher rate), with the acceptance rate for the male applicants unchanged (Table 5 in the Appendix). Bias disclosure with explanations also results in a significant increase in FPR and a significant decrease in FNR, leading to an overall insignificant change in accuracy (Figure 7). Even despite a higher acceptance rate for female applicants, the increase in gender parity is not significant (Figure 7), likely due to the normalizing effect of the acceptance rate of male applicants, which also increases insignificantly.
@@
In the case of indirect bias through a proxy feature, we find that disclosures with explanations have a positive impact on fairness both with respect to perceptions and decision-making. Similar to the without explanations case (§6.2), disclosing both model bias and the association between gender and university while including explanations significantly decreases perceived model fairness (Figure 7). For fairness rating, this effect is significant even without the correlation disclosure. Further, disclosing model bias alone, as well as disclosing model bias along with the correlation between protected and proxy feature with explanations leads to a significant increase in gender parity (Figure 7). This stems from a significantly higher acceptance rate for female applicants., while the acceptance rate for male applicants remains unchanged (Table 5 in the Appendix). This also results in a higher FPR (significant) and a lower FNR (not significant), with an overall drop in accuracy (significant).
In sum, in the case of direct bias, even though bias disclosure with explanations does not improve recognition of model unfairness significantly (possibly because fairness ratings are low even pre-disclosures), it does reduce agreement with the model’s biased decisions, leading to a significantly higher acceptance rate for female applicants (however, decision-making fairness does not improve, possibly because of the normalizing effect of acceptance rate of male applicants, which also increases, albeit insignificantly). This is in stark contrast with the effect of explanations alone (§6.1), which improved recognition of unfairness but led to more biased decisions overall.
Further, in the case of indirect bias, disclosing the model bias and correlations between protected and proxy feature with explanations significantly increases both recognition of unfairness and gender parity in decision-making. We also observe an approximately $1\%$ drop in accuracy in the case of indirect bias after disclosing model bias and correlations with explanations (which might be acceptable in certain cases). Overall, we conclude that neither explanations nor bias or correlation disclosures alone are sufficient. We observe better decision-making fairness outcomes when participants are not only shown the model explanations but also made aware of the biases underlying them.
@@ 6.4. Effect of Joint Intervention
We have seen that explanations alone decrease decision-making fairness, while disclosures with explanations can, in the case of indirect bias, have the opposite effect. Here, we consider the effect of both adding explanations and giving disclosures over including neither (i.e., comparing phase 1 without explanations to phase 2 with explanations).
We find that, for decision-making metrics (accuracy, FPR, and FNR), the effect of joint intervention is never significant (Table 6 in the Appendix). For fairness perception metrics, since the effect of explanations alone and disclosures with explanations pointed in the same direction, as expected, adding both explanations and disclosures also significantly improves recognition of unfairness. In sum, the joint intervention of including both explanations and disclosures (over including neither) helps participants in recognizing model biases but not in correcting them. This is a curious finding as we would expect that giving participants full information about bias in the model and also an explanation of individual prediction would lead to more fair decisions. We believe that this indicates that even though disclosures help undo some of the over-reliance on the model’s biased decisions stemming from the inclusion of explanations, users still tend to be more accepting of model decisions with explanations than without.
@@ 7. Qualitative Results
As described in §4.1, for a selected set of applicants in each phase, participants were asked to write a free text justification for why they agreed or disagreed with AI (or marked it as “neutral”) after marking their agreement and confidence. In addition to encouraging careful thinking, this also helps us gauge the kinds of reasoning participants employ in their decision-making.
As the main goal of our study is to understand how humans interact with AI decisions when the AI is biased, we primarily focus our qualitative analysis on rationales concerning biases. To analyze how participants perceive and use (or discard) the biased feature (gender or university), we consider justifications that directly reference the protected (gender) or proxy feature (university) by using a set of keywords for both. We started with an initial keyword set (e.g., “gender”, “female”, “university”) and, based on reading a subset of the justifications, expanded to include spelling variations (e.g., “skool” and “collage”) and other topically relevant words (e.g., abbreviated names of schools). We discuss our qualitative findings on justifications involving the protected feature in §7.1 and justifications involving the proxy feature in §7.2. Lastly, we discuss additional observations indicating over-reliance based on a random sample of justifications in §7.3.